OpenStudy (anonymous):

x = 3y - 6 2x - 4y = 8 .Solve the system of equations using substitution.

OpenStudy (anonymous):

x=3y-6 2(3y-6)-4y=8 6y-12-4y=8 2y-12=8 2y=20 y=10 x=3*10-6=30-6=24